Dates: September 19 - October 30, 2001
People on site:
Wes Traub (WT), Pete Schloerb (FPS)
Ettore Pedretti (EP), Rafael Millan-Gabet (RMG), John Monnier (JDM), Jean-Philippe Berger (JPB)
Marc Lacasse (MGL), Angela Ahearn (AMA)
Continue work toward implementation of 3-telescope beam combination.
Much of the time was spent recovering from deadly (to the equipment) summer lightning strikes.
In Oct. 1-7, JDM and JPB also tested the visible assymetric coupler.
Status of FPGA/PICNIC Readout
Re-starting LD/Linux Box Ethernet after a Reboot
Notes on CB-Com Connections
Updated Pinout of Anorad/VME Cable
Short Delay Pressure Settings
Ideas on Problems with LD/Linux Drivers
Excerpts from RMG Notes
Log of FPGA/PICNIC Readout Codes (rbf files)